# Largest eigenvalue of positive mean Gaussian matrices

### v1

@@ 1. Introduction
Wigner matrices with independent Gaussian entries (real or complex) have long been a foundation of random matrix theory. The joint distribution of the eigenvalues of such random matrices can be explicitly written down, allowing many interesting properties like the empirical distribution, the largest eigenvalue, gaps between eigenvalues, and local laws to be derived from the closed-form expression. An extension that has gathered considerable attention is the generalization of these results under minimal assumptions on the distribution of the entries, while still maintaining their independence. Such results fall under the category of universality theorems in random matrix theory. For further details on these classical results, the reader is referred to [2], and to [12] for recent developments in local limit laws and their universality. There has also been various works on relaxing the independence assumptions, especially when considering the empirical spectral distribution of the matrices ([3, 4, 8, 1, 9]). Additionally, there have been recent developments concerning bounds on the spectral norm of Wigner matrices with dependent entries ([11, 15]). These works consider general entries with summability conditions on the correlations. It is well known that in the independent case with zero-mean entries, and under some moment conditions, the largest eigenvalue of the matrix scaled by $\sqrt{N}$ converges to $2$ almost surely and, after appropriate scaling and centering, converges weakly to the Tracy-Widom law.
@@ 2. The set-up
The goal is to study the asymptotic behaviour of the largest eigenvalue of $A_{N}$, which will be denoted by $\lambda_{1}(A_{N})$, after a centring and suitable assumptions on $(Z_{ij})$. Interestingly, $\lambda_{1}(A_{N})$ will require no scaling in the regime considered below.
@@
(2.2) $\sum_{i=-\infty}^{\infty}\sum_{j=-\infty}^{\infty}|R(i,j)|<\infty\,,$
@@
(2.3) $\sum_{i=-\infty}^{\infty}\sum_{j=-\infty}^{\infty}R(i,j)\neq 0\,.$
@@
$\sigma^{2}=4\sum_{i=-\infty}^{\infty}\sum_{j=-\infty}^{\infty}R(i,j)\,.$
@@ 3. The proof
An immediate implication of the above is that for all $v\in S^{N-1}$, there exists $x\in S(\varepsilon,N)$ such that $\|v-x\|\leq\varepsilon$. The Cauchy-Schwartz inequality implies
$|v^{\prime}Wv-x^{\prime}Wx|$ $\leq|\langle v-x,Wv\rangle|+|\langle v-x,Wx\rangle|$
$\leq\|Wv\|\|v-x\|+\|Wx\|\|v-x\|$
$\leq 2\varepsilon\|W\|\,,$
the last line following from the definition of the operator norm and that $\|v-x\|\leq\varepsilon$. This, in conjunction with (3.2), shows
$\|W\|\leq(1-2\varepsilon)^{-1}\max_{x\in S(\varepsilon,N)}|x^{\prime}Wx|.$
@@
${\rm Var}\left(x^{\prime}Wx\right)$ $=4{\rm Var}\left(\sum_{i\in{\mathbb{Z}}}\sum_{j\in{\mathbb{Z}}}x_{i}x_{j}X_{ij}\right)$
$=4\sum_{i\in{\mathbb{Z}}}\sum_{j\in{\mathbb{Z}}}\sum_{k\in{\mathbb{Z}}}\sum_{l\in{\mathbb{Z}}}x_{i}x_{j}x_{k}x_{l}R(i-k,j-l)$
$\leq 4\sum_{i\in{\mathbb{Z}}}\sum_{j\in{\mathbb{Z}}}\sum_{k\in{\mathbb{Z}}}\sum_{l\in{\mathbb{Z}}}\left|x_{i}x_{j}x_{k}x_{l}R(i-k,j-l)\right|$
$=4\sum_{u\in{\mathbb{Z}}}\sum_{v\in{\mathbb{Z}}}|R(u,v)|\sum_{i\in{\mathbb{Z}}}\left|x_{i}x_{i+u}\right|\sum_{j\in{\mathbb{Z}}}\left|x_{j}x_{j+v}\right|$
$\leq 4\sum_{u\in{\mathbb{Z}}}\sum_{v\in{\mathbb{Z}}}|R(u,v)|\,,$
@@
$0<\delta\leq\frac{1}{16}\left(\sum_{u\in{\mathbb{Z}}}\sum_{v\in{\mathbb{Z}}}|R(u,v)|\right)^{-1}\,,$
@@
${\rm Var}\left({\bf 1}^{\prime}W{\bf 1}\right)\sim N^{2}\sigma^{2}\,.$
@@
${\rm Var}\left({\bf 1}^{\prime}W{\bf 1}\right)$
$=4{\rm Var}\left(\sum_{i=1}^{N}\sum_{j=1}^{N}X_{ij}\right)$
$=4\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{N}\sum_{l=1}^{N}{\rm Cov}(X_{ij},X_{k,l})$
(3.6) $=4\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{N}\sum_{l=1}^{N}R(i-k,j-l)$
$=4\sum_{u-1-N}^{N-1}\sum_{v=1-N}^{N-1}R(u,v)$
@@
(3.7) $\left|N^{-2}{\rm Var}\left({\bf 1}^{\prime}W{\bf 1}\right)-\sigma^{2}\right|\leq 4\sum_{u\in{\mathbb{Z}}}\sum_{v\in{\mathbb{Z}}}\left(\frac{|u|}{N}\wedge 1\right)\left(\frac{|v|}{N}\wedge 1\right)|R(u,v)|\,.$
@@
${\bf 1}^{\prime}W^{2}{\bf 1}$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{N}W(i,j)W(j,k)$
(3.10) $=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{N}\left(X_{ij}X_{kj}+X_{ji}X_{jk}\right)+2\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{N}X_{ij}X_{jk}\,.$
@@
$\lim_{N\to\infty}N^{-2}\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{N}{\rm E}\left(X_{ij}X_{kj}\right)=\sum_{i=-\infty}^{\infty}R(i,0)\,,$
@@
$\lim_{N\to\infty}N^{-2}\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{N}{\rm E}\left(X_{ji}X_{jk}\right)=\sum_{i=-\infty}^{\infty}R(0,i)\,.$
@@
$\left|\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{N}{\rm E}(X_{ij}X_{jk})\right|$ $=\left|\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{N}R(i-j,j-k)\right|$
$\leq\sum_{j=1}^{N}\sum_{i=-\infty}^{\infty}\sum_{k=-\infty}^{\infty}|R(i-j,j-k)|$
$=N\sum_{u=-\infty}^{\infty}\sum_{v=-\infty}^{\infty}|R(u,v)|\,,$
(3.8) follows.
For showing (3.9), expand the expectation of the square of the first term in the first summation of (3.10) with the help of Wick’s formula ([14, Chapter 1, Theorem 1]) as
${\rm E}\left[\left(\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{N}X_{ij}X_{kj}\right)^{2}\right]$
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{N}\sum_{i^{\prime}=1}^{N}\sum_{j^{\prime}=1}^{N}\sum_{k^{\prime}=1}^{N}{\rm E}\left(X_{ij}X_{kj}X_{i^{\prime}j^{\prime}}X_{k^{\prime}j^{\prime}}\right)$
@@
$\lim_{N\to\infty}N^{-4}{\rm E}\left[\left(\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{N}X_{ij}X_{kj}\right)^{2}\right]=\left(\sum_{i=-\infty}^{\infty}R(i,0)\right)^{2}\,.$
@@
Let $v$ be the eigenvector (normalized to have norm $1$) of $A_{N}$ corresponding to its largest eigenvalue. Thus,
$\lambda_{1}v=A_{N}v=Wv+2\theta({\bf 1}^{\prime}v){\bf 1}\,,$
a consequence of which is
@@
Standard arguments show that
${\bf 1}^{\prime}v\stackrel{{\scriptstyle P}}{{\longrightarrow}}1\,,$
and hence ${\bf 1}^{\prime}v\neq 0$ whp. Thus,
@@
for some $\varepsilon>0$. In other words, (2.2) can possibly be relaxed to something under which a claim similar to that of Theorem 2.1, albeit with a scaling and a different centring, still holds. In this situation, (3.13) will have to be replaced by
@@ Acknowledgement
The research of the second author was supported through NWO Gravitation Grant NETWORKS 024.002.003. The research of the third author is supported by SERB/CRG/2021/006785, from the Science and Engineering Research Board, Government of India.

### v2

@@ 1. Introduction
Wigner matrices with independent Gaussian entries (real or complex) have long been a foundation of random matrix theory. The joint distribution of the eigenvalues of such random matrices can be explicitly written down, allowing many interesting properties like the empirical distribution, the largest eigenvalue, gaps between eigenvalues, and local laws to be derived from the closed-form expression. An extension that has gathered considerable attention is the generalization of these results under minimal assumptions on the distribution of the entries, while still maintaining their independence. Such results fall under the category of universality theorems in random matrix theory. For further details on these classical results, the reader is referred to [2], and to [14] for recent developments in local limit laws and their universality. There has also been various works on relaxing the independence assumptions, especially when considering the empirical spectral distribution of the matrices ([3, 6, 10, 1, 11]). Additionally, there have been recent developments concerning bounds on the spectral norm of Wigner matrices with dependent entries ([13, 18]). These works consider general entries with summability conditions on the correlations. It is well known that in the independent case with zero-mean entries, and under some moment conditions, the largest eigenvalue of the matrix scaled by $\sqrt{N}$ converges to $2$ almost surely [4] and, after appropriate scaling and centering, converges weakly to the Tracy-Widom law [16].
@@ 2. The set-up
The goal is to study the asymptotic behaviour of the largest eigenvalue of $A_{N}$, which will be denoted by $\lambda_{1}(A_{N})$, after a centering and suitable assumptions on $(Z_{ij})$. Interestingly, $\lambda_{1}(A_{N})$ will require no scaling in the regime considered below.
@@
(2.2) $\sum_{i,j=-\infty}^{\infty}|R(i,j)|<\infty\,,$
@@
(2.3) $\sum_{i,j=-\infty}^{\infty}R(i,j)\neq 0\,.$
@@
$\sigma^{2}=4\sum_{i,j=-\infty}^{\infty}R(i,j)\,.$
@@ 3. The proof
Standard arguments show that
$\|W\|\leq(1-2\varepsilon)^{-1}\max_{x\in S(\varepsilon,N)}|x^{\prime}Wx|;$
see [19, Exercise 4.4.3(b)], for example, for the details.
@@
${\rm Var}\left(x^{\prime}Wx\right)$ $=4{\rm Var}\left(\sum_{i,j\in{\mathbb{Z}}}x_{i}x_{j}X_{ij}\right)$
$=4\sum_{i,j,k,l\in{\mathbb{Z}}}x_{i}x_{j}x_{k}x_{l}R(i-k,j-l)$
$\leq 4\sum_{i,j,k,l\in{\mathbb{Z}}}\left|x_{i}x_{j}x_{k}x_{l}R(i-k,j-l)\right|$
$=4\sum_{u,v\in{\mathbb{Z}}}|R(u,v)|\sum_{i\in{\mathbb{Z}}}\left|x_{i}x_{i+u}\right|\sum_{j\in{\mathbb{Z}}}\left|x_{j}x_{j+v}\right|$
$\leq 4\sum_{u,v\in{\mathbb{Z}}}|R(u,v)|\,,$
@@
$0<\delta\leq\frac{1}{16}\left(\sum_{u,v\in{\mathbb{Z}}}|R(u,v)|\right)^{-1}\,,$
@@
$N^{-2}{\rm Var}\left({\bf 1}^{\prime}W{\bf 1}\right)\to\sigma^{2}\,.$
@@
${\rm Var}\left({\bf 1}^{\prime}W{\bf 1}\right)$ $=4{\rm Var}\left(\sum_{i,j=1}^{N}X_{ij}\right)$
$=4\sum_{i,j,k,l=1}^{N}{\rm Cov}(X_{ij},X_{k,l})$
(3.6) $=4\sum_{i,j,k,l=1}^{N}R(i-k,j-l)$
$=4\sum_{u=1-N}^{N-1}\sum_{v=1-N}^{N-1}R(u,v)$
@@
(3.7) $\left|N^{-2}{\rm Var}\left({\bf 1}^{\prime}W{\bf 1}\right)-\sigma^{2}\right|\leq 4\sum_{u,v\in{\mathbb{Z}}}\left(\frac{|u|}{N}\wedge 1\right)\left(\frac{|v|}{N}\wedge 1\right)|R(u,v)|\,.$
@@
${\bf 1}^{\prime}W^{2}{\bf 1}$ $=\sum_{i,j,k=1}^{N}W(i,j)W(j,k)$
(3.10) $=\sum_{i,j,k=1}^{N}\left(X_{ij}X_{kj}+X_{ji}X_{jk}\right)+2\sum_{i,j,k=1}^{N}X_{ij}X_{jk}\,.$
@@
$\lim_{N\to\infty}N^{-2}\sum_{i,j,k=1}^{N}{\rm E}\left(X_{ij}X_{kj}\right)=\sum_{i=-\infty}^{\infty}R(i,0)\,,$
@@
$\lim_{N\to\infty}N^{-2}\sum_{i,j,k=1}^{N}{\rm E}\left(X_{ji}X_{jk}\right)=\sum_{i=-\infty}^{\infty}R(0,i)\,.$
@@
$\left|\sum_{i,j,k=1}^{N}{\rm E}(X_{ij}X_{jk})\right|$ $=\left|\sum_{i,j,k=1}^{N}R(i-j,j-k)\right|$
$\leq\sum_{j=1}^{N}\sum_{i,k=-\infty}^{\infty}|R(i-j,j-k)|$
$=N\sum_{u,v=-\infty}^{\infty}|R(u,v)|\,,$
and hence (3.8) follows.
For showing (3.9), expand the expectation of the square of the first term in the first summation of (3.10) with the help of Wick’s formula ([17, Chapter 1, Theorem 1]) as
${\rm E}\left[\left(\sum_{i,j,k=1}^{N}X_{ij}X_{kj}\right)^{2}\right]$
$=\sum_{i,j,k=1}^{N}\sum_{i^{\prime},j^{\prime},k^{\prime}=1}^{N}{\rm E}\left(X_{ij}X_{kj}X_{i^{\prime}j^{\prime}}X_{k^{\prime}j^{\prime}}\right)$
@@
$\lim_{N\to\infty}N^{-4}{\rm E}\left[\left(\sum_{i,j,k=1}^{N}X_{ij}X_{kj}\right)^{2}\right]=\left(\sum_{i=-\infty}^{\infty}R(i,0)\right)^{2}\,.$
@@
Let $v$ be an eigenvector (normalized to have norm $1$) of $A_{N}$ corresponding to its largest eigenvalue, satisfying
(3.12) ${\bf 1}^{\prime}v\geq 0\,.$
Thus,
(3.13) $\lambda_{1}v=A_{N}v=Wv+2\theta({\bf 1}^{\prime}v){\bf 1}\,.$
Pre-multiply both sides by $v^{\prime}$ and use the fact that $v$ is a unit-norm vector to get
$\lambda_{1}=v^{\prime}Wv+2\theta({\bf 1}^{\prime}v)^{2}\,.$
Dividing throughout by $N$ and using (3.11) in conjunction with Lemma 3.1 yield
$N^{-1}2\theta({\bf 1}^{\prime}v)^{2}\stackrel{{\scriptstyle P}}{{\longrightarrow}}2\theta.$
Divide both sides by $2\theta$ and take the positive square root to conclude
(3.14) $N^{-1/2}({\bf 1}^{\prime}v)\stackrel{{\scriptstyle P}}{{\longrightarrow}}1,$
(3.12) ensuring that ${\bf 1}^{\prime}v$ appears in the left hand side when the positive root is taken.
A consequence of (3.13) is
@@
It follows from (3.14) that ${\bf 1}^{\prime}v\neq 0$ whp. Thus,
@@
for some $\varepsilon>0$. In other words, (2.2) can possibly be relaxed to something under which a claim similar to that of Theorem 2.1, albeit with a scaling and a different centering, still holds. In this situation, (3.16) will have to be replaced by
@@
Such scenarios, and the behavior of the largest eigenvalue under more general correlation structures, are explored in a recent preprint [5].
@@ Acknowledgement
The authors thank an anonymous referee for a careful reading of the manuscript and suggesting changes that have helped to improved it. The research of the second author was supported through NWO Gravitation Grant NETWORKS 024.002.003. The research of the third author is supported by SERB/CRG/2021/006785, from the Science and Engineering Research Board, Government of India.